# Entropic characterization of Tunneling and State Pairing in a Quasi-Exactly Solvable Sextic Potential

### v1

@@ Abstract
We analyze the (de)localization properties of a quasi-exactly solvable (QES) sextic potential $V_{\text{QES}}(x)=\tfrac{1}{2}(x^{6}+2x^{4}-2(2\lambda+1)x^{2}),$ as a function of the tunable parameter $\lambda\in[-\tfrac{3}{4},6]$. For $\lambda>-\tfrac{1}{2}$, the potential exhibits a symmetric double-well structure, with tunneling emerging for the ground state level at $\lambda\approx 0.732953$. For the lowest energy states $n=0,1,2,3$, we construct physically meaningful variational wavefunctions that $i)$ respect parity symmetry under the transformation $x\rightarrow-x$, $ii)$ exhibit the correct asymptotic behavior at large distances, and $iii)$ allow for exact analytical Fourier transforms. Variational energies match Lagrange Mesh and available exact analytical QES results with relative errors $\simeq 10^{-8}$ for $n=0,1,2$ and $\simeq 10^{-6}$ for the third excited state $n=3$. We demonstrate that entropic measures (Shannon entropy, Kullback–Leibler, and Cumulative Residual Jeffreys divergences) surpass conventional variance-based methods in revealing tunneling transitions, wavefunction symmetry breaking, and quantum state pairing. Our results confirm that the Beckner–Białynicki-Birula–Mycielski entropic uncertainty relation holds across all examined values of $n$ and $\lambda$. The quality of the trial function is also validated by the small $\sim 10^{-10}$ Cumulative Residual Jeffreys divergences from the exact QES solutions.
@@ III.1 Variational results
For $\lambda$ a positive integer number, the finite number of $(\lambda+1)$ exact QES eigenfunctions appear in the form Turbiner1988QuasiexactlysolvablePA
@@
where $P_{n}(x)$ is a polynomial of degree $n=0,1,\lambda$, c.f. (4). The QES solutions do not possess exact analytical Fourier transforms.
@@ III.2 Energy spectrum
Figure 2: QES potential $V^{\rm QES}$: $E$ vs $\lambda$ for the lowest states $n=0,1,2,3$. The horizontal line $E=0$ corresponds to the peak of the central barrier when $\lambda>-\frac{1}{2}$, and to the minimum of the potential (3) when $\lambda<-\frac{1}{2}$, see Fig. 1. For each fixed quantum number $n$, the critical value $\lambda_{c}^{n}$ (vertical dashed lines) denotes the coupling strength at which the $n$-th energy level vanishes, i.e., $E_{n}=0$. It marks the onset of tunneling for the $n$-th quantum state. As $\lambda$ grows, the emergence of level pairing is clearly evident. The vertical line at $\lambda=0$ separates the algebraic and non-algebraic sector of the QES system.
@@ IV.1 Position space
Figure 4: Probability densities $|\psi_{n}(x)|^{2}$ for selected paired states ($n=0,1$ and $n=2,3$) at fixed coupling values $\lambda=0$ (top row), $\lambda=1\quad\text{and}\quad 3.5$ (middle row), and $\lambda=4\quad\text{and}\quad 6$ (bottom row). Left column: low-lying states; right column: higher excited states. As the potential evolves into a double well, each pair forms a symmetric–antisymmetric doublet with growing localization and suppressed tunneling, a signature of quantum state pairing in double-well systems.
@@
At very strong coupling ($\lambda=6$, bottom right panel), the effect is even more dramatic: the higher excited states $\psi_{2}(x)$ and $\psi_{3}(x)$ are nearly identical in envelope and peak positions, with only their parity distinguishing them. The suppression of the wavefunction at the origin becomes greatly pronounced in the symmetric states $n=0,2$, and the localization within the individual wells is nearly total. This behavior signifies that tunneling between wells is now extremely rare, and the paired energy levels are effectively quasi-degenerate.
@@
Figure 5: The QES potential $V^{\rm QES}$ (3) marked by the black curve, and the lowest energy levels $n=0,1,2,3$ at different $\lambda$. As $\lambda$ increases, the potential evolves from a single to a double well, leading to tunneling-induced level splitting, wavefunction penetration into the barrier, and eventual symmetry breaking with localized doublets.
@@ IV.2 Momentum space
Figure 6: Momentum-space probability densities $|\phi_{n}(p)|^{2}$ for eigenstates $n=0,1,2,3$ at different coupling values $\lambda$. For $n=0$, the distribution is symmetric and peaked at $p=0$, consistent with the even and nodeless ground state in position space. Odd-$n$ states ($n=1,3$) exhibit nodes at $p=0$, reflecting the antisymmetry of their real-space wavefunctions. Interestingly, the even state $n=2$ shows a local minimum, not a maximum, at $p=0$, due to its nodal structure in position space. As $\lambda$ increases, all states develop multiple side lobes in momentum space, signaling enhanced localization and tunneling suppression in the double-well regime.
@@
Figure 7: Momentum-space probability densities $|\phi_{n}(p)|^{2}$ for eigenstates $n=0,1$ (left column) and $n=2,3$ (right column), evaluated at fixed coupling strengths $\lambda=0$, $2.5$, $3.5$, and $6$. As the excitation number $n$ increases, the distributions broaden and develop oscillatory features due to increased spatial confinement in the double-well potential. Odd-$n$ states exhibit nodes at $p=0$, while even-$n$ excited states display local minima, reflecting the parity of the position-space wavefunctions. No quasi-degenerate structure arises in momentum space.
@@ V Heisenberg and Shannon Entropic uncertainties
Figure 8: Position uncertainty $\Delta x$ (left panel) and Shannon entropy $S_{x}$ (right panel) as functions of the parameter $\lambda$, for eigenstates $n=0,1,2,3$. While $\Delta x$ increases smoothly with $\lambda$ (except for $n=2$), $S_{x}$ exhibits non-monotonic behavior with distinct local maxima near the critical values $\lambda_{c}^{(n)}$, where the corresponding energy levels cross zero. Notably, $S_{x}$ reveals an additional extremum beyond the set of $\lambda_{c}^{(n)}$, signaling finer structural transitions in the wavefunction, such as partial delocalization, tunneling reshaping, or symmetry breaking, that are not captured by the variance alone. Results were obtained from the optimal trial function (4).
@@
Figure 9: Momentum uncertainty $\Delta p$ (left panel) and Shannon entropy $S_{p}$ (right panel) as functions of the parameter $\lambda$, for eigenstates $n=0,1,2,3$. Both measures exhibit qualitatively similar behavior: an initial decrease with $\lambda$, followed by a rise beyond the critical values $\lambda_{c}^{(n)}$, reflecting wavefunction compression and subsequent fragmentation in momentum space. This coherence between $\Delta p$ and $S_{p}$ indicates that both capture similar aspects of momentum-space structure under the influence of the evolving potential. Unlike the position-space case, no curve crossings occur in $\Delta p$, and the ordering with respect to $n$ is preserved throughout.
@@
It is worth noting that the behavior of $\Delta p$ and $S_{p}$ is qualitatively similar, whereas $\Delta x$ and $S_{x}$ do not exhibit such correspondence, indicating that momentum-space measures capture certain structural features more coherently.
@@ V.1 Heisenberg and entropic inequalities
Figure 10: Heisenberg uncertainty product $\Delta x\,\Delta p$ (left panel) and entropic sum $S_{t}=S_{x}+S_{p}$ (right panel) as functions of the parameter $\lambda$, for eigenstates $n=0,1,2,3$. In the left panel, the dashed red line marks the lower bound $\Delta x\,\Delta p=1/2$, corresponding to the ground state of the harmonic oscillator. The Heisenberg product exhibits a quasi-pairing structure, $(n=0,1)$ and $(n=2,3)$, reminiscent of harmonic oscillator behavior, with mild separation and smooth evolution. In contrast, the entropic sum $S_{t}$, bounded below by $1+\log\pi\approx 2.144$ the dashed red line (right panel), breaks this apparent symmetry: it displays clear distinctions between all four states, revealing enhanced sensitivity to tunneling, state localization, and symmetry breaking near the critical values $\lambda_{c}^{(n)}$. Odd-parity states ($n=1,3$) are less affected by variations in $\lambda$, due to their vanishing amplitude at the center of the potential. For each fixed quantum number $n$, the critical value $\lambda_{c}^{n}$ (vertical dashed lines) denotes the coupling strength at which the $n$-th energy level vanishes, i.e., $E_{n}=0$. The vertical line at $\lambda=0$ separates the algebraic and non-algebraic sector of the QES system.
Figure 10 displays the Heisenberg uncertainty product $\Delta x\,\Delta p$ (left panel) and the total Shannon entropy $S_{t}=S_{x}+S_{p}$ (right panel) as functions of the parameter $\lambda$ for the first four eigenstates ($n=0,1,2,3$) of the sextic QES potential.
@@
Interestingly, for the low lying states $n=0,1,2,3$ the harmonic-like behavior in $\Delta x\Delta p$ persists beyond $\lambda=-\tfrac{1}{2}$, where the potential undergoes a qualitative transition to a double-well structure. Although the potential develops a central barrier and symmetry breaking becomes possible, the states maintain a value of $\Delta x\,\Delta p$ effectively equal to that of the harmonic oscillator. This is somewhat surprising, as double-well systems typically induce delocalization and bimodal wavefunction shapes. The persistence of low uncertainty values suggests a highly coherent and symmetric ground state that remains delocalized over both wells before significant tunneling asymmetry sets in.
@@ VI Quantifying Tunneling and Pairing in the QES Potential Using KL and CRJ Divergences
Figure 11: Cumulative Residual Jeffreys Divergence (CRJ) of the eigenstate pairs $(n=0,1)$ (solid red line) and $(n=2,3)$ (dashed blue line) versus $\lambda$.
Figure 11 shows the Cumulative Residual Jeffreys (CRJ) divergence HSR2019 as a function of the parameter $\lambda$, comparing the spatial probability densities of the eigenstate pairs $(n=0,1)$ (solid red line) and $(n=2,3)$ (dashed blue line).
@@
In contrast, the blue curve shows the CRJ divergence for the higher pair $(n=2,3)$, which begins at low values and remains relatively flat in the single-well regime. Around $\lambda\approx 2$, the divergence begins to increase, reaching a maximum near $\lambda\approx 3.2$, and then decreases again. The appearance of this maximum corresponds to the onset of pairing between $n=2$ and $n=3$: prior to this point, the states are structurally distinct, but as the double-well deepens, they begin to form another near-degenerate symmetric-antisymmetric pair. The peak in the CRJ divergence indicates the point of maximal asymmetry in their spatial structures, typically when one state begins to delocalize across both wells while the other remains nodal or partially localized. After the maximum, the decline in CRJ signals increasing similarity as the pairing becomes more complete (see in Figure 4 (d) and (f) ).
Most notably, the curve Kullback-Leibler (KL) divergence (black solid line) decreases monotonically with $\lambda$, approaching zero for $\lambda\gtrsim 3$. This behavior indicates a progressive increase in similarity between the ground and first excited states. In the deep double-well regime, these states form a nearly degenerate pair corresponding to symmetric and antisymmetric superpositions of localized states in each well. As a result, their probability densities (which are symmetric under squaring) become nearly identical, and the KL divergence between them vanishes. This reflects the emergence of tunneling-induced state pairing, a hallmark of double-well systems.
@@
Figure 12: CRJ divergence between QES sextic and harmonic oscillator densities for $n=0,1,2$, as functions of the parameter $\lambda$. Despite similar Heisenberg uncertainty products, the QES states exhibit nontrivial structural deviations from their HO counterparts, especially beyond the critical coupling $\lambda_{c}^{(n=0)}$.
@@
These contrasting behaviors underscore the distinct ways in which QES states reorganize in response to the double-well transition. In general, for $\lambda>1$, while some states regain harmonic-like features due to symmetric delocalization, others increasingly depart due to node formation, asymmetry, or partial localization. A brief summary of the main findings of the present study is presented in Table 1.
@@
Finally, it is worth mentioning that the CRJ divergence (of the order $\sim 10^{-10}$) between the available exact QES probability densities and their variational approximation confirms the accuracy and reliability of our trial function (4), see Table 2.
00 $n=0$ $1.24\times 10^{-11}$
$1$ $n=0$ $2.80\times 10^{-10}$
$n=2$ $9.39\times 10^{-11}$
$2$ $n=0$ $5.38\times 10^{-11}$
$n=2$ $5.84\times 10^{-10}$
$3$ $n=0$ $5.38\times 10^{-11}$
$n=2$ $1.58\times 10^{-9}$
Table 2: Comparison between exact QES probability densities $\rho_{{}_{\rm QES}}$ and the approximate variational densities $\rho_{{}_{\rm trial}}$ obtained from (4) for different values of $\lambda=0,1,2,3$ and lowest even states $n=0,2$.

### v2

@@ Abstract
We analyze the (de)localization properties of a quasi-exactly solvable (QES) sextic potential $V_{\text{QES}}(x)=\tfrac{1}{2}(x^{6}+2x^{4}-2(2\lambda+1)x^{2}),$ as a function of the tunable parameter $\lambda\in[-\tfrac{3}{4},6]$. For $\lambda>-\tfrac{1}{2}$, the potential exhibits a symmetric double-well structure, with tunneling emerging for the ground state level at $\lambda\approx 0.732953$. For the lowest energy states $n=0,1,2,3$, we construct physically meaningful variational wavefunctions that $i)$ respect parity symmetry under the transformation $x\rightarrow-x$, $ii)$ exhibit the correct asymptotic behavior at large distances, and $iii)$ allow for exact analytical Fourier transforms. Variational energies match Lagrange Mesh and available exact analytical QES results with relative errors $\simeq 10^{-8}$ for $n=0,1,2$ and $\simeq 10^{-6}$ for the third excited state $n=3$. We demonstrate that entropic measures (Shannon entropy, Kullback–Leibler, and Cumulative Residual Jeffreys divergences) surpass conventional variance-based methods in revealing tunneling transitions, wavefunction symmetry breaking, and quantum state pairing. Our results confirm that the Beckner–Białynicki-Birula–Mycielski entropic uncertainty relation holds across all examined values of $n$ and $\lambda$. The quality of the trial density functions is also validated by the small $\sim 10^{-10}$ Cumulative Residual Jeffreys divergences from the available exact QES solutions.
@@ III.1 Variational results
For $\lambda$ a positive integer number, the finite number of $(\lambda+1)$ exact QES eigenfunctions appears in the form Turbiner1988QuasiexactlysolvablePA
@@
where $P_{N}(x)$ is a polynomial of degree $N=0,1,2\,\ldots,\lambda$, c.f. (4). The value $N=0$ corresponds to the ground state $(n=0)$, $N=1$ to the second excited state ($n=2$), $N=2$ to the fourth excited state ($n=4$) and so on. The QES solutions do not possess exact analytical Fourier transforms.
@@ III.2 Energy spectrum
Figure 2: QES potential $V^{\rm QES}$: Energy levels $E$ vs. coupling $\lambda$ for the lowest states $n=0,1,2,3$. The line $E=0$ corresponds to the central barrier peak for $\lambda>-\frac{1}{2}$, and to the potential minimum for $\lambda<-\frac{1}{2}$ (see Fig. 1). Vertical dashed lines indicate critical couplings $\lambda_{c}^{n}$ where $E_{n}=0$, signaling tunneling onset. The line at $\lambda=0$ separates algebraic and non-algebraic sectors.
@@ IV.1 Position space
Figure 4: Probability densities $|\psi_{n}(x)|^{2}$ for selected paired states ($n=0,1$ and $n=2,3$) at fixed coupling values $\lambda=0$ (top row), $\lambda=1\quad\text{and}\quad 3.5$ (middle row), and $\lambda=4\quad\text{and}\quad 6$ (bottom row). Left column: low-lying states; right column: higher excited states.
@@
At very strong coupling ($\lambda=6$, bottom right panel), the effect is even more dramatic: the higher excited states $\psi_{2}(x)$ and $\psi_{3}(x)$ are nearly identical in envelope and peak positions, with only their parity distinguishing them. The suppression of the wavefunction at the origin becomes greatly pronounced in the symmetric states $n=0,2$, and the localization within the individual wells is nearly total. This behavior signifies that tunneling between wells is now extremely rare, and the paired energy levels are effectively quasi-degenerate. We note that, in the present study, we do not present the wavefunctions themselves but rather their associated probability densities $|\psi^{2}(x)|$ (and $|\phi^{2}(p)|$), since these are the relevant quantities that enter into the informational entropic analysis.
@@
Figure 5: The QES potential $V^{\rm QES}$ (3) (black curve) and the lowest energy levels $n=0,1,2,3$ at various $\lambda$. As $\lambda$ increases, the potential transitions from a single to a double well, producing level splitting, barrier penetration, and eventual symmetry breaking with localized doublets.
@@ IV.2 Momentum space
Figure 6: Momentum-space densities $|\phi_{n}(p)|^{2}$ for $n=0,1,2,3$ at various $\lambda$. Even states peak or dip at $p=0$, depending on their nodal structure; odd states vanish there due to antisymmetry.
@@
Figure 7: Momentum-space densities $|\phi_{n}(p)|^{2}$ for $n=0,1$ (left) and $n=2,3$ (right) at $\lambda=0,2.5,3.5,6$. As $n$ increases, the distributions broaden and become oscillatory due to stronger spatial confinement. Odd states vanish at $p=0$; even excited states show local minima. No quasi-degeneracy appears in momentum space.
@@ V Heisenberg and Shannon Entropic uncertainties
Figure 8: Position uncertainty $\Delta x$ (left) and Shannon entropy $S_{x}$ (right) vs. $\lambda$ for states $n=0,1,2,3$. While $\Delta x$ grows smoothly (except for $n=2$), $S_{x}$ shows non-monotonic behavior with peaks near $\lambda_{c}^{(n)}$, and for $n=2$ an additional extremum beyond them, indicating structural changes in the wavefunction not captured by variance. Results use the optimal trial function (4).
@@
Figure 9: Momentum uncertainty $\Delta p$ (left) and Shannon entropy $S_{p}$ (right) vs. $\lambda$ for $n=0,1,2,3$. The rise beyond the critical values $\lambda_{c}^{(n)}$, indicate compression and fragmentation in momentum space. They capture similar structural features. In $\Delta p$ no curve crossings appear, and the ordering with respect to $n$ is preserved.
@@
It is worth noting that the behavior of $\Delta p$ and $S_{p}$ is qualitatively similar, whereas $\Delta x$ and $S_{x}$ do not exhibit such correspondence, indicating that momentum-space measures capture certain structural features more clearly.
@@ V.1 Heisenberg and entropic inequalities
Figure 10: Heisenberg product $\Delta x\,\Delta p$ (left) and entropic sum $S_{t}=S_{x}+S_{p}$ (right) as functions of $\lambda$ for $n=0,1,2,3$. Each panel includes its corresponding theoretical lower bound (dashed red line).
Figure 10 illustrates that within the range $-0.75\leq\lambda\leq 0.2$, the Heisenberg product exhibits behavior characteristic of the harmonic oscillator. In contrast, the entropic sum $S_{t}$ deviates from this apparent symmetry, underscoring the greater sensitivity of entropic measures to subtle variations in the underlying potential structure.
@@
Interestingly, for the low-lying states $n=0,1,2,3$ the harmonic-like behavior in $\Delta x\Delta p$ persists beyond $\lambda=-\tfrac{1}{2}$, where the potential undergoes a qualitative transition to a double-well structure. Although the potential develops a central barrier and symmetry breaking becomes possible, the states maintain a value of $\Delta x\,\Delta p$ effectively equal to that of the harmonic oscillator. This is somewhat surprising, as double-well systems typically induce delocalization and bimodal wavefunction shapes. The persistence of low uncertainty values suggests a highly coherent and symmetric ground state that remains delocalized over both wells before significant tunneling asymmetry sets in. We will return to this point later.
@@ VI Quantifying Tunneling and Pairing in the QES Potential Using KL and CRJ Divergences
Figure 11: CRJ Divergence for the eigenstate pairs $(n=0,1)$ (red curve), $(n=2,3)$ (blue curve) and KL Divergence for the eigenstate pairs $(n=0,1)$ (black curve)versus $\lambda$.
Figure 11 shows the Cumulative Residual Jeffreys (CRJ) divergence HSR2019 as a function of the parameter $\lambda$, comparing the spatial probability densities of the eigenstate pairs $(n=0,1)$ (red curve) and $(n=2,3)$ (blue curve).
@@
Notably, the Kullback–Leibler (KL) divergence, represented by the black curve, exhibits a monotonic decrease as a function of $\lambda$, approaching zero for $\lambda\gtrsim 3$. This monotonic trend signals a growing similarity between the ground and first excited states. In the deep double-well regime, this observation is consistent with the CRJ divergence results (see Fig. 11), where both states form a nearly degenerate pair consisting of symmetric and antisymmetric combinations of localized wavefunctions in each well. Since the probability densities—being quadratic in the wavefunction—are nearly identical in this regime, the KL divergence between them also vanishes. This behavior confirms the formation of tunneling-induced state pairs, a hallmark of quantum systems governed by symmetric double-well potentials.
In contrast, the blue curve shows the CRJ divergence for the higher pair $(n=2,3)$, which begins at low values and remains relatively flat in the single-well regime. Around $\lambda\approx{\lambda_{c}^{n=0}}$, the divergence begins to increase, reaching a maximum near $\lambda\approx 3.2$, and then decreases again. The appearance of this maximum corresponds to the onset of pairing between $n=2$ and $n=3$: prior to this point, the states are structurally distinct, but as the double-well deepens, they begin to form another near-degenerate symmetric-antisymmetric pair. The peak in the CRJ divergence indicates the point of maximal asymmetry in their spatial structures, typically when one state begins to delocalize across both wells while the other remains nodal or partially localized. After the maximum, the decline in CRJ signals increasing similarity as the pairing becomes more complete (see in Figure 4 (d) and (f) ).
@@
Figure 12: CRJ divergence between QES sextic and harmonic oscillator densities for $n=0,1,2$ vs $\lambda$. Despite similar Heisenberg uncertainty products, the QES states exhibit nontrivial structural deviations from their HO counterparts.
@@
These contrasting behaviors underscore the distinct ways in which QES states reorganize in response to the double-well transition. In general, for $\lambda>1$, while some states regain harmonic-like features due to symmetric delocalization, others increasingly depart due to node formation, asymmetry, or partial localization. A brief summary of the main findings of the present study is presented in Table 2. Finally, it is worth mentioning that the CRJ divergence (of the order $\sim 10^{-10}$) between the available exact QES probability densities and their variational approximation confirms the accuracy and reliability of our trial function (4), see Table 1.
00 $n=0$ $1.24\times 10^{-11}$
$1$ $n=0$ $2.80\times 10^{-10}$
$n=2$ $9.39\times 10^{-11}$
$2$ $n=0$ $5.38\times 10^{-11}$
$n=2$ $5.84\times 10^{-10}$
$3$ $n=0$ $5.38\times 10^{-11}$
$n=2$ $1.58\times 10^{-9}$
Table 1: Comparison between exact QES probability densities $\rho_{{}_{\rm QES}}$ and the approximate variational densities $\rho_{{}_{\rm trial}}$ obtained from (4) for $\lambda=0,1,2,3$ and lowest even states $n=0,2$.